OpenStudy (anonymous):

Evaluate the composite function f(g(x)) for x = 31.

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

g(31)= 6, f(g(31))=f(6)=9*6+1=55.hope helps.

OpenStudy (anonymous):

thank you!